OpenStudy (anonymous):

Evaluate the following expression using the values given: Find -3a2 - b3 + 3c2 - 2b3 if a = 2, b = -1, and c = 3.

OpenStudy (anonymous):

15

OpenStudy (anonymous):

Proof ?

OpenStudy (anonymous):

-3 x 2 x2 = - 12 -12 - -1 x 3 = -9 -9 + 3 x 3 x2 =9 9 - 2 x -1 x 3 = 15 answer is 15

OpenStudy (anonymous):

It was wrong

OpenStudy (anonymous):

!?!?!?! where did i go wrong?

OpenStudy (anonymous):

I dont know but the answer was wrong

OpenStudy (anonymous):

sorry